You accept a job that pays $12.50 an hour. You are told that after a two month probationary period, your hourly wage will be increased to $15 an hour. What percent raise will you receive after the two month period?
Answer by sofiyac(983)

so you'll be getting 2.50 more per hour
12.50--100%
2.50---x% cross multiply
12.50x=(2.50*100) solve for x, divide both sides by 12.50
x=(2.50*100/12.50)=20%